3. Explain how the level of saving is determined in the simple Keynesian consumption function. What is the effect of an increase in disposable income on the level of saving?
1
Expert's answer
2019-10-02T09:12:50-0400

Keynesian formula for consumption is

Consumption = autonomous consumption + marginal propensity to consume × disposable income.

We can write it as

"\u0421 = a + b*Yd"


In the same time disposable income consist of consumption and savings, so:

"Yd = C +S"


From that two formulas we're getting a formula for savings:

"S = Yd - C"

"S = -a + (1 - b)*Yd"


So −a is the level of autonomous saving and (1 − b) is the marginal propensity to save.


According to Keynesians, with the increase of disposable income - people consume a smaller fraction of their income: the average propensity to consume decreases"(APC = C\/Yd)" . And the average propensity to save increases"(APS = S\/Yd)" .